OpenStudy (anonymous):

Find the slope of the line through the following pairs of points. (5, -1) (10,2) what is the rise? what is the run?

OpenStudy (anonymous):

m=2+1/10-5=-3/5

OpenStudy (anonymous):

slope= rise /run

OpenStudy (anonymous):

\[Find the slope of the line through the following pairs of points. (5, -1)(10,2) Plot each pair of points, draw a line through them. \]